A new car is purchased for 19900 dollars. The value of the car depreciates at 7.5% per year. What will the value of the car be, to the nearest cent, after 8 years?


1
Expert's answer
2022-02-28T12:03:44-0500

"Amount" "=P(1-r\/100)"n

"A=19900(1-(7.5\/100))"8

"A=19900(1-0.075)"8

"A" "=""19900(0.925)"8

"A=19900\times0.5359618307"

"A=10,665.64 dollars"